OpenStudy (anonymous):

Select all that justify the following statement. 10 + x 12 = 12x + 10 commutative - addition distributive associative - multiplication symmetric commutative - multiplication associative - addition identity - addition

OpenStudy (anonymous):

10+ x * 12=12x+10

OpenStudy (anonymous):

commutative - multiplication

OpenStudy (anonymous):

commutative law twice. \[10+x12=10+12x\] that was the commutative law of multiplication. now \[10+12x=12x+10\] is the commutative law of addition. so you used them both